Solve each system using any method.
a. -3x + y = -3 b. b = -4a + 1
4x + 2y = 14 a - 3b = 10


c. For further thought, what is the value of x + y - a + b

Answer by Boreal(15235)

rewrite first as
6x-2y=6, multiplying by -2
4x+2y=14
10x=20
x=2
y=3
(2, 3)
-
substitute in the second
a-3(-4a+1)=10
a+12a-3=10
a=1
b=-3
(1, -3)
x+y-a+b=5-1-3=1
Do they mean (x+y)-(a+b)? That answer is 3